OpenStudy (newarf):

after taking 20% off the original price victor paid $14 what was the original price of the book

OpenStudy (irish_waffl3):

16.80

OpenStudy (mysticfury):

I would set this up as a basic algebra equation. Since $14 is the price after a 20% discount then $14 is is 80% of the original price. $14/80=x/100 (100% would be the original price) 1400=80x or 1400/80 = $17.5! You can double check the work by taking 20% off the $17.5 like this 17.5 x 0.20 = 3.5 17.5 - 3.5 = 14 :)

OpenStudy (mysticfury):

16.80 is incorrect because if we take 20% off of 16.80 we get 13.44 16.80 x 0.20 = 3.36 16.80 - 3.36 = 13.44 not 14$

OpenStudy (newarf):

@mysticfury what's the answer then

OpenStudy (kropot72):

The answer has already been given in the calculation by @mysticfury.

OpenStudy (mysticfury):

^^ yes it has :)

OpenStudy (mysticfury):

but since you arent gonna read it all to know how to do it , it originally cost $17.50

OpenStudy (newarf):

I read it just don't understand it

OpenStudy (mysticfury):

ah okay thats fine, as long as your read it and not just glanced it over for the answer :)

OpenStudy (newarf):

ok